The following data apply to questions 3 through 10. Sebastian Company, which manufactures electrical switches, uses a standard cost system and carries all inventories at standard. The standard manufacturing overhead costs per switch are based on direct labor hours and are shown below: Variable overhead (5 hours @ $12 per direct manufacturing labor hour) $ 60 Fixed overhead (5 hours @ $15* per direct manufacturing labor hour) 75 Total overhead per switch $135 *Based on capacity of 200,000 direct manufacturing labor hours per month. The following information is available for the month of December:  46,000 switches were produced although 40,000 switches were scheduled to be produced.  225,000 direct manufacturing labor hours were worked at a total cost of $5,625,000.  Variable manufacturing overhead costs were $2,750,000.  Fixed manufacturing overhead costs were $3,050,000. 9. What is the manufacturing overhead flexible-budget variance for December a. $10,000 F. b. $40,000 U. c. $50,000 U. d. $100,000 U. 10. The manufacturing overhead spending variance for December was a. $10,000 F. b. $40,000 U. c. $50,000 U. d. $100,000 U.
Answer #1
Flexible Budget
Variable Overhead $   2,760,000 =46000*5*12
Fixed Overhead $   3,000,000 =200000*15
$   5,760,000
Actual
Variable Overhead $   2,750,000
Fixed Overhead $   3,050,000
$   5,800,000


Manufacturing Overhead Flexible Budget Variance = Actual Overhead - Flexible Budget Overhead
= $5800000 - $5760000 = $40000 (U)

Answer is b. $40000 U

Fixed overhead spending variance = Actual Overhead - Budgeted Overhead
= $3050000 - $3000000 = $50000 (U)

Variable Overhead Spending Variance = Actual Overhead - Actual hours x Standard Rate
= $275000 - 225000 x $12 = $50000 (U)

Manufacturing Overhead spending Variance = $50000 + $50000 = $100000 (U)

Answer is d. $100,000 U